# Meta-automatic Sequences

### v1

@@ Abstract
Nested (or meta-Fibonacci) recurrences, such as the recurrence used to define Hofstadter’s $Q$-sequence, along with the digit-based recurrences that underlie automatic sequences are of interest from both number-theoretic and combinatorial points of view. In this direction, Allouche and Shallit showed how the frequency sequence of a variant of the $Q$-sequence is $2$-automatic. This inspires us to introduce what may be seen as a natural combination of the recurrences for meta-Fibonacci and automatic sequences, by introducing the concept of a meta-automatic sequence. We show how it is possible to construct a meta-automatic sequence that is not denestable in terms of not being reducible, in a specific sense formalized in this paper, to certain digit-based recurrences for automatic sequences. This motivates our study of the non-denestable, meta-automatic sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ introduced in this paper. For each of these integer sequences, we prove explicit DFAO evaluations, together with $4$-uniform morphisms, and we also consider the factor complexities of these sequences.
@@ 1 Introduction
Informally, a sequence $(a(n))_{n\geq 0}$ is said to satisfy a nested or meta-Fibonacci recurrence if $a(n)$, for all sufficiently large $n$, can be expressed in terms of $a$ evaluated at an expression involving a value given as a term in $(a(n))_{n\geq 0}$. The classical and prototypical example is Hofstadter’s $Q$-sequence [19, p. 137], which is indexed in the On-Line Encyclopedia of Integer Sequences [28] as A005185, and which satisfies $Q(1)=Q(2)=1$ and
@@
for $n>2$. In contrast to nested recurrences, digit-based recurrences (which may also be referred to as divide-and-conquer recurrences) define $a\big(k^{t}n+j\big)$ (for $j\in\{0,1,\ldots,k^{t}-1\}$) as a function of previously computed values of the form $a\big(k^{t^{\prime}}n^{\prime}+j^{\prime}\big)$ for $t^{\prime}<t$ (and for $j^{\prime}\in\{0,1,\ldots,k^{t^{\prime}}-1\}$). These digit-based recurrences are used to define automatic sequences [3, 9], which are of great importance in many areas of number theory, computer science, and combinatorics.
The depth of the study of meta-Fibonacci sequences is evidenced by the many open problems related to Hofstadter’s $Q$-sequence and generalizations. Notably, to this day, it is not even known whether or not $Q(n)$ is defined for all integers $n>0$. This together with the roles played by automatic sequences within many disciplines in mathematics and computer science lead us to consider a new and interdisciplinary subject that may be thought of as “combining” the recursions for meta-Fibonacci and automatic sequences.
As it turns out, relationships between meta-Fibonacci and automatic sequences have previously been considered. Notably, Allouche and Shallit [4], in 2012, showed that the frequency sequence of a variant of Hofstadter’s $Q$-sequence satisfies a digit-based recurrence, so that this variant sequence is $2$-automatic. In this context, they proved an automaticity criterion [4, Theorem 2.2] reviewed in Section 1.2 below.
In this paper, we introduce the concept of a meta-automatic sequence, which, informally, can be thought of as extending both the Allouche–Shallit criterion and the recursions for meta-Fibonacci sequences as in (1). For example, in contrast to a rule such as $a(4n+2)=a(2n)$ used to define an automatic sequence, we consider rules such as
@@
where the argument of the right-hand side depends on the value $a(n)$. Recurrences as in (2) do not, in general, fall under the Allouche–Shallit automaticity criterion [4, Theorem 2.2], since the argument on the right-hand side of (2) is not necessarily equivalent to a polynomial in $n$ of the desired form.
The idea of combining the recursions associated with automatic and meta-Fibonacci sequences is motivated by many past references concerning meta-Fibonacci sequences, in addition to Allouche and Shallit’s work on the automaticity of the frequency sequence for Hofstadter’s $Q$-sequence [4]. In this direction, we highlight notable references on the Hofstadter $Q$-sequence [16, 17, 18, 25], and on variants and generalizations of the Hofstadter $Q$-sequence [1, 2, 5, 12, 14, 30] and on meta-Fibonacci sequences [7, 8, 10, 11, 13, 15, 20, 21, 22, 26, 29].
@@ 1.1 Outline
A formal definition of the term meta-automatic sequence is given in Section 2 below.
As we demonstrate, it is possible for the recurrence for a meta-automatic sequence to, in a specific sense, degenerate in such a way so that it reduces to a digit-based recurrence. To formalize what should be meant by such reductions, we introduce the concept of denestability. As it turns out, there exist meta-automatic sequences that are not denestable, and we introduce the first construction of a sequence of this form. In this direction, Theorem 23 shows that the integer sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ that we introduce in this paper are not denestable. This leads to the following question motivating the topic of our paper: Under what conditions do meta-automatic recurrences produce automatic sequences?
The binary sequences $(a(n))_{n\geq 0}$ under consideration in this paper are mainly balanced in the sense that $a(2n)+a(2n+1)=1$. This condition provides a defining property of the Thue–Morse sequence, and so the meta-automatic sequences introduced in this paper may be seen as “meta-Thue–Morse” sequences.
We prove that in base $2$, meta-automatic sequences that are binary and balanced are ultimately periodic (Lemma 14). However, as we later clarify, by passing to a base-$4$ digit template that preserves a dyadic balance, this can be used to construct balanced, binary, and meta-automatic sequences that are not ultimately periodic. In this direction, a key idea is given by how the balance constraint linearizes each meta-automatic selector via an $\mathbb{F}_{2}$-affine operation (Lemma 10). This motivates our study of the following meta-automatic sequences:
  • Meta(1) $\mathcal{M}_{1}$ (A392736): a mixed template in which one of the two branch rules uses a meta-automatic selector. The XOR linearization yields a $4$-state DFAO (Section 3).
  • Meta(2) $\mathcal{M}_{2}$ (A391614): a “pure” two-selector template in which both branch rules are meta-automatic. The linearized recurrence has an affine structure over $\mathbb{F}_{2}^{2}$ governed by the matrix $\bigl(\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\bigr)$, and the sequence admits the factorization $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, where $q$ is an explicit bit-masking operator and $\mathbf{t}$ denotes the Thue–Morse sequence (Theorem 35).
Remarkably, despite $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ sharing the same automaton size of $4$ states, the two sequences differ sharply in structural complexity. Notably, we find that $\mathcal{M}_{2}$ admits the evaluation $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, while $\mathcal{M}_{1}$ requires a lifted size-two kernel. Despite this shared automaton size, their factor-complexity functions, studied in Section 5, differ markedly.
For each of the above sequences, we give an MSB-first, base-$4$ DFAO, an explicit $4$-uniform morphism, and factor-complexity results. The broader theme is that the interplay between balancedness and $\mathbb{F}_{2}$-linearity can bring meta-automatic, digit-based recurrences into the realm of automatic sequences, even when criteria for automaticity, such as the above referenced Allouche–Shallit criterion [4], do not apply.
@@ 1.2 Preliminaries
For background on automatic sequences, we refer to the appropriate texts from Allouche and Shallit [3] and from Shallit [27]. We briefly recall some definitions that are useful or neeed, for our purposes. In this direction, the concept of a $k$-kernel [3, §6.6] is required in our work.
@@
For $k\geq 2$, a $k$-automatic sequence may then be defined as a sequence $U=(U(n))_{n\geq 0}$ such that $K_{k}(U)$ is finite. Equivalently, this sequence is $k$-automatic if there exists a DFAO that, on input the base-$k$ representation of $n$, outputs $U(n)$, again referring to Allouche and Shallit’s text for background.
Definition 2.
Fix a base $k\geq 2$. By a digit-based recurrence we mean a system of recurrences that defines $a(kn+r)$ for digits $r\in\{0,1,\ldots,k-1\}$ in terms of values of $a(m)$ at strictly smaller indices $m<kn+r$ (typically among $m=n,2n,\dots$).
Let $\mathbf{t}:\mathbb{N}_{0}\to\{0,1\}$ denote the Thue–Morse sequence, defined by $\mathbf{t}(0)=0$ and
@@
Equivalently, $\mathbf{t}(n)=s_{2}(n)\bmod 2$, where $s_{2}(n)$ is the sum of the binary digits of $n$.
Since much of our work is closely related to an automaticity criterion due to Allouche and Shallit [4], it is necessary, for our purposes, to review this criterion, as below.
@@
for $n\geq 0$. Moreover, for integers $q\geq 2$ and $t\geq 1$, we write $s(q,t)=\frac{1-q^{t+1}}{1-q}$ in place of the number of pairs $(i,j)$ of integers $i$ and $j$ such that $0\leq i\leq t$ and $0\leq j\leq q^{i}-1$. We then order expressions of the form $U_{q,i,j}$ by writing $U_{1}=U$, $U_{2}$, $\ldots$, $U_{s(q,t)}$. We also let $m(q,t)=q^{t+1}-1$.
@@
(Allouche &\& Shallit, 2012) Let the terms of $(U(n))_{n\geq 0}$ be in a finite set $\mathcal{A}$, and let $q\geq 2$ denote an integer. Then $U$ is $q$-automatic if there are nonnegative integers $t$, $a$, $b$, and $n_{0}$ together with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions from $\mathcal{A}^{a+b+s(q,t)}$ to $\mathcal{A}$ such that
@@ 2 Meta-automatic sequences
Let $q>0$ and $i_{0}>0$ and $j_{0}\geq 0$ denote integers. For some $v\geq 0$, and for indices $k\in\{1,2,\ldots,v\}$, let $\mathcal{S}_{k}$ and $i_{k}>0$ and $j_{k}\geq 0$ be integers, subject to the condition that $i_{k}<i_{0}$ for all indices $k$. By analogy with (3), and with the assumption that the codomain of $U$ is contained in $\mathbb{N}_{0}$, we write
$U_{q,i_{0},j_{0},\mathcal{S}_{1},i_{1},j_{1},\ldots,\mathcal{S}_{v},i_{v},j_{v}}=U\Big(q^{i_{0}}n_{0}+j_{0}+\mathcal{S}_{1}U(q^{i_{1}}n_{1}+j_{1})+\cdots+\mathcal{S}_{v}U(q^{i_{v}}n_{v}+j_{v})\Big).$
@@
Let the terms of $(U(n))_{n\geq 0}$ be in a finite subset $\mathcal{A}$ of $\mathbb{N}_{0}$. Let $q\geq 2$ be an integer. Then $U$ is $q$-meta-automatic if there are nonnegative integers $t$, $a$, $b$, and $n_{0}$ along with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions to $\mathcal{A}$ (with domains given below) such that, for all $n\geq n_{0}$ and for all $j\in\{0,1,\ldots,m(q,t)\}$, the value $U\big(q^{t+1}n+j\big)$ is equal to $f_{j}$ evaluated at a tuple of the form
$\Big(U^{-a}(n),\ldots,U^{-1}(n),U^{0}(n),U^{1}(n),\ldots,U^{b}(n),\\ U_{q,i_{0}^{(1)},j_{0}^{(1)},\mathcal{S}_{1}^{(1)},i_{1}^{(1)},j_{1}^{(1)},\ldots,\mathcal{S}_{v_{1}}^{(1)},i_{v_{1}}^{(1)},j_{v_{1}}^{(1)}},\ldots,U_{q,i_{0}^{(w)},j_{0}^{(w)},\mathcal{S}_{1}^{(w)},i_{1}^{(w)},j_{1}^{(w)},\ldots,\mathcal{S}_{v_{w}}^{(w)},i_{v_{w}}^{(w)},j_{v_{w}}^{(w)}}\Big)$
for some $w\geq 0$ and for integers $i_{k_{1}}^{(k_{2})}\leq t$ and $j_{k_{1}}^{(k_{2})}$ and $\mathcal{S}_{k_{1}}^{(k_{2})}$, and where the domain of $f_{j}$ is $\mathcal{A}^{a+b+w+1}$.
@@
Let the integer sequence $(\mathcal{T}(n))_{n\geq 0}$ be defined so that $\mathcal{T}(0)=0$ and so that
@@
According to Definition 4, the sequence $(\mathcal{T}(n))_{n\geq 0}$ is 22-meta-automatic. It can be shown that this sequence agrees, for $n\geq 1$, with the OEIS sequence A039982, which is defined in a different way and without the use of meta-recursions.
Definition 6.
For a $q$-meta-automatic sequence $(U(n))_{n\geq 0}$, adopting notation from Definition 4, a selector refers to an expression given by $f_{j}$ evaluated at the tuple displayed in Definition 4, i.e., with such an expression being equal to $U(q^{t+1}n+j)$.
Example 7.
For the $2$-meta-automatic sequence in Example 5, the expression $\mathcal{T}(n-\mathcal{T}(n-1))$ is a selector.
Binary sequences of the following form are a main object of study in our work.
Definition 8.
A binary sequence $a:\mathbb{N}_{0}\to\{0,1\}$ is said to be balanced if
$a(2n)+a(2n+1)=1\qquad(n\geq 0).$ (4)
Example 9.
The Thue–Morse sequence may be seen as a prototypical instance of a balanced (binary) sequence.
As in the following result, we let $\oplus$ denote the additive operation in the field $\mathbb{F}_{2}$.
@@
This follows in a direct way from (4). ∎
Remark 11.
The selector a(2n+1a(n))a(2n+1-a(n)) and the XOR expression $a(2n+1)\oplus a(n)$ compute the same function of the triple $(a(n),a(2n),a(2n+1))$. Both read the same two independent bits $(a(n),a(2n))$, since balancedness gives a(2n+1)=1a(2n)a(2n+1)=1-a(2n). The balance constraint does not reduce the number of value dependencies, but it ensures that the dependency is $\mathbb{F}_{2}$-affine: The selector, viewed as a Boolean function, is addition in $\mathbb{F}_{2}$.
2.1 Denestability
We introduce two notions that distinguish the sequences studied in this paper from those covered by the automaticity criterion of Allouche and Shallit [4, Theorem 2.2].
Definition 12.
A meta-automatic recurrence for a sequence $(U(n))_{n\geq 0}$ is denestable if every selector associated with an expression of the form $U\big(q^{t+1}n+j\big)$ can, for sufficiently large $n\geq n_{0}$, be written as an expression of the form given in the right-hand side of the identity in Theorem 3, without increasing the size of kernel.
Example 13.
The integer sequence $(\mathcal{T}(n))_{n\geq 0}$ defined in Example 5 is denestable, since the given recursion is such that $\mathcal{T}(2n)=1$ for all $n\geq 1$, providing, together with the relation $\mathcal{T}(2n+1)=1-\mathcal{T}(n)$, a digit-based recurrence for $\mathcal{T}$.
A weaker notion, relative to Definition 12, is that of $\mathbb{F}_{2}$-linearizability, which, informally, allows the balance constraint to convert meta-automatic selectors into $\mathbb{F}_{2}$-affine maps at fixed arguments. This is later clarified.
In base $2$, the most direct way to introduce a meta-automatic argument while keeping the binary balancedness (4) is, informally, to try to define the even subsequence using a “one-level” selector at the $n$-scale. For instance, set

### v2

@@ Abstract
Nested (or meta-Fibonacci) recurrences, such as the recurrence used to define Hofstadter’s $Q$-sequence, along with the digit-based recurrences that underlie automatic sequences are of interest from both number-theoretic and combinatorial points of view. In this direction, Allouche and Shallit showed how the frequency sequence of a variant of the $Q$-sequence is $2$-automatic. This inspires us to introduce what may be seen as a natural combination of the recurrences for meta-Fibonacci and automatic sequences, by introducing the concept of a meta-automatic sequence. We exhibit two binary meta-automatic sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ whose defining recurrences do not satisfy the Allouche–Shallit automaticity criterion directly, and this is formalized in our paper. For each of these integer sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, we prove explicit DFAO evaluations, together with 4-uniform morphisms, and we also consider the factor complexities of these sequences.
@@ 1 Introduction
A nested recurrence is a recurrence in which a term $a(n)$ is defined using values of $a$ at indices that themselves depend on earlier values of the sequence. The classical example is Hofstadter’s $Q$-sequence [20, p. 137], indexed in the On-Line Encyclopedia of Integer Sequences [30] as A005185, defined by $Q(1)=Q(2)=1$ and
@@
for $n>2$. Recurrences as in (1) are also known as meta-Fibonacci recurrences. Other examples of integer sequences satisfying nested recurrences include the sequence A244477 studied by Golomb [18]. Many problems concerning Hofstadter’s $Q$-sequence and its generalizations remain open to this day. Notably, it is not known whether or not $Q(n)$ is even defined for all $n>0$.
In contrast to nested recurrences, digit-based recurrences (also called divide-and-conquer recurrences) define $a(k^{t}n+j)$ for $j\in\{0,1,\ldots,k^{t}-1\}$ as a function of previously computed values $a(k^{t^{\prime}}n^{\prime}+j^{\prime})$ with $t^{\prime}<t$ and $j^{\prime}\in\{0,1,\ldots,k^{t^{\prime}}-1\}$. Digit-based recurrences can be used to define automatic sequences [3, 9], which appear in many areas of number theory, computer science, and combinatorics.
Allouche and Shallit [4] showed that the frequency sequence of a variant of Hofstadter’s $Q$-sequence satisfies a digit-based recurrence, and, moreover, is $2$-automatic. They also proved an automaticity criterion [4, Theorem 2.2] reviewed in Section 1.2 below. This inspires us to introduce the concept of a meta-automatic sequence, which may, informally, be thought of as “combining” the concepts of meta-Fibonacci and automatic sequences. This is formalized below. For example, in contrast to a rule such as $a(4n+2)=a(2n)$ used to define an automatic sequence, we consider rules such as
@@
in which the argument of $a$ on the right-hand side depends on the value $a(n)$. A recurrence of the form suggested in (2) would not, in general, satisfy the Allouche–Shallit automaticity criterion, since the argument on the right is not of the required polynomial form.
The idea of combining automatic and meta-Fibonacci recurrences is motivated by past work on meta-Fibonacci sequences, in addition to the above referenced work by Allouche and Shallit [4]. We point to references on the Hofstadter $Q$-sequence [16, 17, 19, 27], on variants and generalizations of the Hofstadter $Q$-sequence [1, 2, 5, 12, 14, 32], and on meta-Fibonacci sequences [7, 8, 10, 11, 13, 15, 21, 22, 23, 28, 31].
@@ 1.1 Outline
A formal definition of a meta-automatic sequence is given in Section 2. A meta-automatic recurrence may, after simplification, reduce to a digit-based recurrence, in which case the Allouche–Shallit criterion applies directly. We formalize, in this paper, how meta-automatic sequences are not necessarily reducible in this way.
A binary sequence $a:\mathbb{N}_{0}\to\{0,1\}$ is said to be balanced if
$a(2n)+a(2n+1)=1\qquad(n\geq 0).$ (3)
The binary sequences under consideration in this paper are, for the most part, balanced. In a related way, much of this paper makes use of the additive operation underlying the field $\mathbb{F}_{2}=\{0,1\}$, letting $\oplus$ denote this operation, which may be referred to as XOR, with $0\oplus 0=1\oplus 1=0$ and $0\oplus 1=1\oplus 0=1$. For balanced binary sequences, Lemma 5 gives the identity
$a(2n+1-a(n))=a(2n+1)\oplus a(n).$
We prove that balanced, binary, meta-automatic sequences are, for the cases covered in Lemma 6, ultimately periodic. To obtain non-periodic examples, we therefore work with base-$4$ recurrences. We consider:
  • The integer sequence $\mathcal{M}_{1}$ (A392736), defined via a base-$4$ recurrence in which one of the rules in the recurrence is nested in the sense outlined above. After applying Lemma 5 below, the pair $(\mathcal{M}_{1}(n),b(n))$, for $b(n)$ defined in (8) below, yields a 44-state DFAO (Section 3).
  • The integer sequence $\mathcal{M}_{2}$ (A391614), defined via a base-$4$ recurrence in which both of the rules in the recurrence are nested in the sense outlined above. After applying Lemma 5, the pair $\bigl(\mathcal{M}_{2}(n),\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n))\bigr)$ obeys an affine rule over $\mathbb{F}_{2}^{2}$ with linear part $\bigl(\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\bigr)$. The sequence satisfies $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, where $q$ is an explicit bit-masking operator and $\mathbf{t}$ is the Thue–Morse sequence (Theorem 26).
Although $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ both have $4$-state DFAOs, we have that $\mathcal{M}_{2}$ admits the evaluation $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, whereas $\mathcal{M}_{1}$ is described by means of the pair $(\mathcal{M}_{1}(n),b(n))$. Their factor-complexity functions, studied in Section 5, also differ. For each of the sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, we give a base-$4$ DFAO (MSB-first) and an explicit $4$-uniform morphism, and we also investigate the factor complexity functions for $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in Section 5.
@@ 1.2 Preliminaries
For background on automatic sequences, we refer to the texts of Allouche and Shallit [3] and of Shallit [29]. We proceed with the following definition, which may be used, as below, to define automatic sequences.
@@
For $k\geq 2$, a $k$-automatic sequence is a sequence $U=(U(n))_{n\geq 0}$ such that $K_{k}(U)$ is finite. Equivalently, the sequence $U$ is $k$-automatic if there exists a DFAO such that, upon inputting the base-$k$ representation of $n$, this DFAO outputs $U(n)$. See Allouche and Shallit’s text [3] for details. A prototypical instance of an automatic sequence is the Thue–Morse sequence $\mathbf{t}:\mathbb{N}_{0}\to\{0,1\}$ defined so that $\mathbf{t}(0)=0$ and
@@
Equivalently, $\mathbf{t}(n)=s_{2}(n)\bmod 2$, where $s_{2}(n)$ is the sum of the binary digits of $n$, noting that the balanced condition in (3) holds. We introduce, in this paper, a variant of an automaticity criterion due to Allouche and Shallit [4], which we now recall.
@@
for $n\geq 0$. Moreover, for integers $q\geq 2$ and $t\geq 1$, we write $s(q,t)=\frac{1-q^{t+1}}{1-q}$ in place of the number of pairs $(i,j)$ of integers $i$ and $j$ such that $0\leq i\leq t$ and $0\leq j\leq q^{i}-1$. Fix an ordering of these pairs $(i,j)$, with $(0,0)$ appearing first, and write the corresponding sequences as $U_{1}=U,\,U_{2},\ldots,\,U_{s(q,t)}$. The choice of ordering plays no role in what follows. We also let $m(q,t)=q^{t+1}-1$.
@@
(Allouche and Shallit, 2012) Let the terms of $(U(n))_{n\geq 0}$ be in a finite set $\mathcal{A}$, and let $q\geq 2$ denote an integer. Then $U$ is $q$-automatic if there are nonnegative integers $t$, $a$, $b$, and $n_{0}$ together with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions from $\mathcal{A}^{a+b+s(q,t)}$ to $\mathcal{A}$ such that
@@ 2 Meta-automatic sequences
Let $q>0$, $i_{0}>0$, and $j_{0}\geq 0$ be integers. For some $v\geq 0$, and for indices $k\in\{1,2,\ldots,v\}$, let $c_{k}$, $i_{k}>0$, and $j_{k}\geq 0$ be integers, with $i_{k}<i_{0}$ for all $k$. By analogy with (4), and assuming that the codomain of $U$ is contained in $\mathbb{N}_{0}$, we write
$U_{q,i_{0},j_{0},c_{1},i_{1},j_{1},\ldots,c_{v},i_{v},j_{v}}(n)=U\Big(q^{i_{0}}n+j_{0}+c_{1}U(q^{i_{1}}n+j_{1})+\cdots+c_{v}U(q^{i_{v}}n+j_{v})\Big).$
The following definition may be thought of as formalizing the intuition behind recurrences as in (2).
@@
Suppose that each term of $(U(n))_{n\geq 0}$ is in a finite subset $\mathcal{A}$ of $\mathbb{N}_{0}$, and let $q\geq 2$ be an integer. The sequence $U$ is $q$-meta-automatic if there exist nonnegative integers $t$, $a$, $b$, $n_{0}$ together with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions to $\mathcal{A}$, with domains as below, such that, for all $n\geq n_{0}$ and all $j\in\{0,1,\ldots,m(q,t)\}$, the value $U(q^{t+1}n+j)$ equals $f_{j}$ evaluated at the tuple
$\Big(U^{-a}(n),\,\ldots,\,U^{-1}(n),\,U^{0}(n),\,U^{1}(n),\,\ldots,\,U^{b}(n),\\ U_{q,i_{0}^{(1)},j_{0}^{(1)},c_{1}^{(1)},i_{1}^{(1)},j_{1}^{(1)},\ldots,c_{v_{1}}^{(1)},i_{v_{1}}^{(1)},j_{v_{1}}^{(1)}}(n),\,\ldots,\\ U_{q,i_{0}^{(w)},j_{0}^{(w)},c_{1}^{(w)},i_{1}^{(w)},j_{1}^{(w)},\ldots,c_{v_{w}}^{(w)},i_{v_{w}}^{(w)},j_{v_{w}}^{(w)}}(n)\Big)$
for some $w\geq 0$ and integers $i_{k_{1}}^{(k_{2})}\leq t$, $j_{k_{1}}^{(k_{2})}$, and $c_{k_{1}}^{(k_{2})}$, where the domain of $f_{j}$ is $\mathcal{A}^{a+b+w+1}$.
@@
Define $(\mathcal{T}(n))_{n\geq 0}$ by $\mathcal{T}(0)=0$ and
@@
Then $(\mathcal{T}(n))_{n\geq 0}$ is $2$-meta-automatic in the sense of Definition 3. For $n\geq 1$, this sequence agrees with the OEIS sequence A039982, which was defined without nested recurrences.
@@
The first identity follows from (3) by case analysis on $a(n)\in\{0,1\}$, and similarly for the second. ∎
2.1 Base-4 recurrences
We now record a result that motivates the use of base-$4$ recurrences in the sequel. We let